OpenStudy (anonymous):

Please help: x^3-9x^3/2+8=0

OpenStudy (anonymous):

put \(u=x^{\frac{3}{2}}\) and solve \[u^2-9u+8=0\] the go back and replace \(u\) by \(x^{\frac{3}{2}}\)

OpenStudy (anonymous):

I get as far as (u-8)(u-1)=0 then u=8 and u=1 but what do I do after that?